# Unveiling the Nature and Fate of the Almost-Dark Cloud AGC 226178 through Mapping

### v1

@@ 3.2 Small-scale morphology and kinematics
While the FAST observations detect faint and extended gas structures, with poor resolving power on galactic scales, VLA observations can revolve gas in individual objects, albeit at relatively worse column density sensitivity. Fig.4 compares the integrated spectra of AGC 226178 from ALFALFA, FAST and the VLA. The overall velocity profiles of ALFALFA and FAST are in broad agreement with each other. The VLA observations appear to miss the faint, low-velocity tail shown in the single-dish spectra. Particularly, the natural-weighted image recovers in total, which is of the ALFALFA flux. Given the analysis in the previous section, the missing flux of the VLA data is expected to be mostly associated with extended features (resulting from tidal or ram pressure stripping) beyond the main body of AGC 226178.
Fig. 5 presents the VLA moment 0 maps produced with robust (top-left; bottom-left), natural (top-right; bottom-middle) and -taperred natural (bottom-right) weightings. Similar to that discovered by Cannon et al. (2015) based on a D-configuration map, AGC 226178 appears highly centrally concentrated and single-peaked in our lowest resolution map with -taperred weighting. However, the distribution is resolved into two major overdensities, approximately aligned with the brightest blue stellar clusters along the north-east to south-west direction in the optical main body of AGC 226178. The peak of the central clump has N() = , whereas the peak of the clump in the south-west has N() = . In addition to the two major peaks aligned along the north-east to south-west direction, the north-east to south-west direction across the central peak is also stretched, matching the spatial locations of the relatively faint blue stellar clusters to the north-west and south-east end (Jones et al. 2022a). In the robust-weighted map, which has a linear resolution of 0.6 kpc, the major clump in the natural-weighted map is further resolved into several sub-peaks indicating a hierarchical gas structure in AGC 226178.
@@
The whole cloud exhibits a complex morphology, its optical counterpart appears elongated. It also shows significant extension perpendicular to the main axis. While a velocity gradient is present, its origin is unclear and cannot be conclusively attributed to rotation. The inclination angle is also highly uncertain given the limited spatial resolution. Given the overall ragged yet round-ish morphology of the intensity distribution, we adopt a series of circular apertures to derive the radial profile. Circularly averaged radial profile of the column density N() extracted from the natural-weighted moment 0 map is shown in Fig. 6. The measured central N() reaches 1020.8 cm-2 ( 5 pc-2). The N() radial profile can be fitted with a Gaussian function. The fitting returns a FWHM of 2.7 kpc. Thus, by subtracting the mean FWHM of the beam in quadrature, the intrinsic FWHM is estimated to be 2.4 kpc. The deconvolved central N is cm-2 ( 6 pc-2).
@@
We modelled the system with a rotational velocity field using the 3D tilted ring fitter 3DBarolo
, which is well-suited for modelling a velocity field that is significantly affected by beam smearing (Di Teodoro & Fraternali 2015). Given the complexity of the velocity field of AGC 226178 (Fig. A1) and the tentative rotation-like gradient discussed above, we inspected each channel of the natural-weighted image cube and manually masked features that are not associated with the central elongated structure. After the masking step, there remains 0.12 Jy , accounting for of the total flux in the natural-weighted cubes and corresponding to a mass of .
@@
To estimate the dynamical mass of the main body (i.e., excluding the low-velocity tail structure), we adopt the following equation from Hoffman et al. (1996):
where is the rotational velocity of the system, is the velocity dispersion, in km s-1, whereas is the radius of the system in kpc.
@@ 3.4 Resolved clumps in AGC 226178
The VLA observations reveal that the distribution in AGC 226178 is highly fragmented and clumpy. To quantify the properties of the clumps, we used the CLUMPFIND
algorithm in the ‘Starlink’ package CUPID
333https://starlink.eao.hawaii.edu/starlink/CUPID to detect clumps on the naturally weighted data cube.
Although the higher-resolution robust-weighted cube appears to reveal more structures, only the relatively bright ones can be robustly detected and measured, due to the poorer sensitivity. So here we focus on the natural-weighted image cube. As input for running CLUMPFIND
, an RMS noise level of 0.4 , measured from the natural-weighted cube, is used. Pixels below 1 RMS are excluded from detection, each clump must have a spatial area greater than the beam area, and adjacent peaks are considered disjoint only if the dip between them exceeds 0.2 RMS and their separation is greater than 2 pixels. To illustrate the clumpy structures, the channel maps are presented in Fig. 8.
A total of 23 clumps were detected by CUPID
. After visual inspection, we only kept the brightest 10 clumps which have an integrated signal-to-noise ratio larger than . The relevant properties of these clumps are listed in Table 2. The reported spatial size of each clump has been corrected for beam-smearing by deconvolving the instrumental beam width, assuming a Gaussian profile, and the velocity width has been similarly corrected for the velocity resolution of the data. Distribution of the clumps in the 3D space of RA, DEC, and velocity is illustrated in Fig. 9, where the left and middle panels are for two different viewing angles. Each clump is plotted as an ellipsoid, of which the length of the three principle axes is, proportional to the spatial size and velocity width, and the ellipsoid is colour-coded according to the integrated flux of each clump. The right panel shows a 3D visualisation of the image cube created with Slicer
444https://www.slicer.org/. We also present the intensity map and velocity profiles of these clumps in Fig. 10.
@@
We point out that, as with previous similar studies in the literature, the search for clumps is strongly limited by the resolving power of the data. This is because the structure of the interstellar medium is hierarchical in nature, as demonstrated by the finer details revealed in our robust-weighted maps (e.g., Fig. 5). Moreover, the size of the clumps (Table 2) reaches up to kpc scales which is an order of magnitude larger than the largest gravitationally-bound giant molecular clouds. Therefore, the clumps identified here represent the finest overdensities in the distribution that can be resolved by our VLA data.
@@ 3.5 Relation between gas and star formation
We use the H emission from star-forming regions as a tracer of the current star formation in AGC 226178. AGC 226178 has previously been observed using Very Large Telescope (VLT)/Multi Unit Spectroscopic Explorer (MUSE, Bacon et al. 2014), as part of the observing programme 0101.B-0376A (P.I: R. Muñoz), the results have been presented in Bellazzini et al. (2022). The MUSE 1′ 1′ field for AGC 226178 (also known as BC3 in Bellazzini et al. (2022)), was centred at RA, DEC = (191.677299, 10.36919), observed on 28 Feb 2019, with a seeing disc . regions were identified by Bellazzini et al. (2022) using a narrow-band image created by integrating a 5 spectral window centred on the H emission line. Regions with peak intensities exceeding 3 above the background were selected and H fluxes were measured within circular apertures of radius . We make direct use of the regions identified and characterised in their analysis; the details of the identification and measurement procedures can be found in their paper. The greyscale background in Fig. 10 displays continuum-subtracted H line emission image from MUSE, integrated over the wavelength range 6591.54–6605.29Å. The plot illustrates the projected spatial relationship between the regions and clumps.
@@
We associate the clumps identified in Section 3.4 with the regions based on their proximity in both spatial location and velocity. It turns out that 4 clumps (ID: 1, 2, 4, 5) have associated regions, which are indicated in Table 2. The SFR for the whole system and each individual region was calculated from the line flux (corrected for extinction using the Balmer decrement technique) of the regions, employing the calibration provided by Kennicutt (1998), which assumes a Salpeter initial mass function (IMF). However, we note that this –SFR conversion may be uncertain in our case, as the assumption of constant (stationary) star formation over timescales of several Myr may not hold. Moreover, at such low SFR levels (especially for ), stochastic sampling of the IMF may lead to an underestimation of the total star formation (see, e.g., Boselli et al. 2009, 2018a).
@@
In Fig. 11, we explore the relation between the SFR and mass of the clumps. Literature samples of nearby galaxies and AGC 226178 as a whole are also plotted for comparison. There is a correlation between the mass and SFR, but with substantial scatter. The clumps associated with star formation (blue squares in Fig. 11) in AGC 226178 generally follow the relation observed for nearby galaxies (left panel). The -based star formation efficiency SFR/ of AGC 226178 and its clumps fall at the lower end of the range spanned by nearby galaxies (right panel). This broad consistency of AGC 226178 with normal galaxies supports the view that the -SFR relation does not reflect a causal connection, but rather is a manifestation of star formation being self-regulated by heating (partly contributed by star formation) and radiative cooling of the interstellar medium (e.g., Taylor & Webster 2005).
@@
In Fig. 12, we further explore the structural properties of the clumps and their connection to star formation. The mass is plotted against clump size in the left panel, and the velocity dispersion of the clumps is plotted against size in the right panel. Mass-size relations corresponding to several characteristic surface densities are also plotted for comparison. Particularly, an surface density of is about the observed maximum on (sub-)kpc scales of nearby galaxies (e.g., Wong & Blitz 2002). It is thought that at higher surface densities is mostly transformed to molecular hydrogen. A surface density of is about the theoretical critical value for a gas cloud (in protogalactic discs) that can be self-shielded from cosmic background radiation (Taylor & Webster 2005).
@@
In the left panel of Fig. 12, we also overplot the five almost dark clouds (including AGC 226178) from Cannon et al. (2015). There is a tight mass-size correlation for these sources, similar to the well-known tight correlation observed for ordinary galaxies (e.g., Broeils & Rhee 1997; Wang et al. 2016). The sense of the correlation is broadly consistent with nearly constant average surface densities in different systems. The existence of a similarly tight mass-size correlation for clumps and almost dark clouds suggests a self-regulated structural property of interstellar neutral gas, regardless of environment. With that being said, we note that all of the sources plotted in Fig. 12 lie below the line of maximum surface density mentioned above, suggesting that these clouds or clumps are -dominated. In addition, the star formation-free clumps have systematically lower surface densities than the clumps with star formation and other almost dark clouds. The average surface densities of the star formation-free clumps are systematically lower than the theoretical mentioned above. The intracluster radiation field (ignored in the theoretical calculation) is generally much stronger than the cosmic background radiation field. It is therefore quite possible that the clumps without associated star formation are not sufficiently self-shielded, which may inhibit the formation of molecular hydrogen and, consequently, suppress star formation.
@@
In the clump size-velocity dispersion plot shown in the right panel of Fig. 12, there is a significant correlation (the pearson correlation coefficient = 0.86), which is expected for a hierarchical fragmentation governed by turbulence. The best-fit relation is:
@@
The logarithmic slope of 6.5 is much steeper than that ( 0.4-0.5; e.g. Larson 1981; Heyer et al. 2009) of the Larson relations for molecular clouds. A cloud dominated by self-gravity has a logarithmic slope close to 0.5, by assuming a constant surface density. Therefore, our clumps are not expected to be bound by gravity. In addition, we note that the velocity dispersion of clumps associated with star formation is systematically (for given size) larger than those not associated with star formation. This suggests a significant feedback of star formation on the neighbouring gas.
ID | (deg) | (deg) | (km s-1) | (mJy km s-1) | (arcsecarcsec) | (km s-1) | |
1 | 191.67907 | 10.36797 | 1581.2 | 58.7 | 12.818.1 | 7.2 | BC3s9, BC3s10, BC3s12, BC3s13, BC3s14 |
2 | 191.67497 | 10.36574 | 1591.1 | 56.9 | 16.918.1 | 9.0 | BC3s15, BC3s18, BC3s23 |
3 | 191.67847 | 10.37008 | 1570.1 | 29.7 | 16.412.9 | 6.2 | – |
4 | 191.67606 | 10.37303 | 1577.3 | 18.8 | 10.811.3 | 5.0 | BC3s4, BC3s24W, BC3s26W |
5 | 191.68145 | 10.36332 | 1573.7 | 12.7 | 7.410.1 | 4.0 | BC3s19 |
6 | 191.67465 | 10.36463 | 1578.1 | 13.5 | 8.111.9 | 4.0 | BC3s20 |
7 | 191.67906 | 10.36520 | 1568.9 | 18.1 | 14.310.9 | 3.2 | – |
8 | 191.68031 | 10.36443 | 1591.1 | 4.4 | 5.75.9 | 0.3 | – |
9 | 191.67911 | 10.37473 | 1574.3 | 8.1 | 9.58.3 | 2.6 | – |
10 | 191.68112 | 10.36895 | 1589.5 | 8.3 | 6.411.2 | 1.5 | – |
@@ 4 Discussion
4.1 The connection of AGC 226178 with nearby galaxies
@@
Jones et al. (2022a) uncovered a faint, low signal-to-noise bridge between AGC 226178 and VCC 2034, based on the ALFALFA observation. If the bridge were confirmed, it would be a strong evidence for a physical connection between the two systems. Based on this discovery, Jones et al. (2022a) came to the conclusion that AGC 226178 was formed from a gas stripped from VCC 2034, The unusually high gas-phase metallicity ( 0.5 ) of AGC 226178 indicates that its gas was probably from a galaxy with a much higher stellar mass ( 108.4 ; Jones et al. 2022a) rather than self-enriched. The model from Junais et al. (2022) showed that VCC 2034 may have experienced a ram-pressure stripping event starting 150 Myr ago, losing about of gas, which could comfortably account for the formation of AGC 226178.
@@
Besides the neutral gas, ionised gas emission may also be used to probe the environmental stripping event (Jones et al. 2022b). According to the radiative hydrodynamical simulations of Lee et al. (2022), besides the relatively bright H emitting star-forming regions, there exists diffuse ionised gas in the ram pressure stripped tails. Diffuse H emission from the stripped tails may be detected at surface brightness levels fainter than 6 1038 erg s-1 kpc-2. This is within the reach of the Virgo Environmental Survey Tracing Ionised Gas Emission (VESTIGE) survey (Boselli et al. 2018b), whose 1 surface brightness sensitivity is 3.4 1038 erg s-1 kpc-2 for a distance of 16.5 Mpc (Jones et al. 2022b). However, as discussed in detail by Boselli et al. (2022), this ionised phase may be short-lived. In RPS-affected galaxies, gas is likely stripped as neutral hydrogen and then becomes ionised and eventually heated when mixing with the intracluster medium (ICM). Consequently, diffuse ionised gas is only observed in a handful of late-type galaxies in Virgo, most of which are currently undergoing active ram pressure stripping (Longobardi et al. 2020; Sardaneta et al. 2022). Nevertheless, given the signs of potential environmental interaction in AGC 226178, we examined the VESTIGE narrow-band H image, we did not find any diffuse emission in between and beyond the optical main body of AGC 226178 and VCC 2034 at the current epoch.
@@
4.2 The nature of AGC 226178: A stripped, disintegrating gas cloud or a mini-halo?

### v2

@@ 3.2 Small-scale morphology and kinematics
While the FAST observations detect faint and extended gas structures, with poor resolving power on galactic scales, VLA observations can revolve gas in individual objects, albeit at relatively worse column density sensitivity. Fig.4 compares the integrated spectra of AGC 226178 from ALFALFA, FAST and the VLA. The overall velocity profiles of ALFALFA and FAST are in broad agreement with each other. The VLA observations appear to miss the faint, low-velocity tail shown in the single-dish spectra. Particularly, the natural-weighted image recovers in total, which is of the ALFALFA flux. Given the analysis in the previous section, the missing flux of the VLA data is expected to be mostly associated with extended features (resulting from tidal or ram pressure stripping) beyond the main body of AGC 226178. Fig. 5 presents the VLA moment 0 maps produced with robust (top left; bottom left), natural (top right; bottom middle) and -tapered natural (bottom right) weightings.
Similar to that discovered by Cannon et al. (2015) based on a D-configuration map, AGC 226178 appears highly centrally concentrated and single-peaked in our lowest resolution map with -taperred weighting. However, the distribution is resolved into two major overdensities, approximately aligned with the brightest blue stellar clusters along the north-east to south-west direction in the optical main body of AGC 226178. The peak of the central clump has N() = , whereas the peak of the clump in the south-west has N() = . In addition to the two major peaks aligned along the north-east to south-west direction, the north-east to south-west direction across the central peak is also stretched, matching the spatial locations of the relatively faint blue stellar clusters to the north-west and south-east end (Jones et al. 2022a). In the robust-weighted map, which has a linear resolution of 0.6 kpc, the major clump in the natural-weighted map is further resolved into several sub-peaks indicating a hierarchical gas structure in AGC 226178.
@@
The whole cloud exhibits a complex morphology, its optical counterpart appears elongated. It also shows significant extension perpendicular to the main axis. While a velocity gradient is present, its origin is unclear and cannot be conclusively attributed to rotation. The inclination angle is also highly uncertain given the limited spatial resolution. Given the overall ragged yet round-ish morphology of the intensity distribution, we adopt a series of circular apertures to derive the radial profile. Circularly averaged radial profile of the column density N() extracted from the natural-weighted moment 0 map is shown in Fig. 6. The measured central N() reaches 1020.8 cm-2 ( 5 pc-2). The N() radial profile can be fitted with a Gaussian function. The fitting returns a full width at half maximum (FWHM) of 2.7 kpc. Thus, by subtracting the mean FWHM of the beam in quadrature, the intrinsic FWHM is estimated to be 2.4 kpc. The deconvolved central N is cm-2 ( 6 pc-2).
@@
We modelled the system with a rotational velocity field using the 3D tilted ring fitter 3DBarolo, which is well-suited for modelling a velocity field that is significantly affected by beam smearing (Di Teodoro & Fraternali 2015). Given the complexity of the velocity field of AGC 226178 (Fig. 14) and the tentative rotation-like gradient discussed above, we inspected each channel of the natural-weighted image cube and manually masked features that are not associated with the central elongated structure. After the masking step, there remains 0.12 Jy , accounting for of the total flux in the natural-weighted cubes and corresponding to a mass of .
@@
To estimate the dynamical mass of the main body (i.e. excluding the low-velocity tail structure), we adopt the following equation from Hoffman et al. (1996):
where is the rotational velocity of the system, is the velocity dispersion, in km s-1, whereas is the radius of the system in kiloparsec.
@@ 3.4 Resolved clumps in AGC 226178
The VLA observations reveal that the distribution in AGC 226178 is highly fragmented and clumpy. To quantify the properties of the clumps, we used the CLUMPFIND algorithm in the Starlink package CUPID333https://starlink.eao.hawaii.edu/starlink/CUPID to detect clumps on the naturally weighted data cube.
Although the higher-resolution robust-weighted cube appears to reveal more structures, only the relatively bright ones can be robustly detected and measured, due to the poorer sensitivity. So here we focus on the natural-weighted image cube. As input for running CLUMPFIND, an RMS noise level of 0.4 , measured from the natural-weighted cube, is used. Pixels below 1 RMS are excluded from detection, each clump must have a spatial area greater than the beam area, and adjacent peaks are considered disjoint only if the dip between them exceeds 0.2 RMS and their separation is greater than 2 pixels. To illustrate the clumpy structures, the channel maps are presented in Fig. 8.
A total of 23 clumps were detected by CUPID. After visual inspection, we only kept the brightest 10 clumps which have an integrated signal-to-noise ratio larger than . The relevant properties of these clumps are listed in Table 2. The reported spatial size of each clump has been corrected for beam-smearing by deconvolving the instrumental beam width, assuming a Gaussian profile, and the velocity width has been similarly corrected for the velocity resolution of the data. Distribution of the clumps in the 3D space of RA, DEC, and velocity is illustrated in Fig. 9, where the left and middle panels are for two different viewing angles. Each clump is plotted as an ellipsoid, of which the length of the three principle axes is, proportional to the spatial size and velocity width, and the ellipsoid is colour-coded according to the integrated flux of each clump. The right panel shows a 3D visualization of the image cube created with Slicer.444https://www.slicer.org/ We also present the intensity map and velocity profiles of these clumps in Fig. 10.
@@
We point out that, as with previous similar studies in the literature, the search for clumps is strongly limited by the resolving power of the data. This is because the structure of the interstellar medium is hierarchical in nature, as demonstrated by the finer details revealed in our robust-weighted maps (e.g. Fig. 5). Moreover, the size of the clumps (Table 2) reaches up to kiloparsec scales which is an order of magnitude larger than the largest gravitationally bound giant molecular clouds. Therefore, the clumps identified here represent the finest overdensities in the distribution that can be resolved by our VLA data.
@@ 3.5 Relation between gas and star formation
We use the H emission from star-forming regions as a tracer of the current star formation in AGC 226178. AGC 226178 has previously been observed using Very Large Telescope (VLT)/Multi Unit Spectroscopic Explorer (MUSE, Bacon et al. 2014), as part of the observing programme 0101.B-0376A (P.I: R. Muñoz), the results have been presented in Bellazzini et al. (2022). The MUSE 1′ 1′ field for AGC 226178 (also known as BC3 in Bellazzini et al. (2022)), was centred at RA, DEC = (191.677299, 10.36919), observed on 28 Feb 2019, with a seeing disc . regions were identified by Bellazzini et al. (2022) using a narrow-band image created by integrating a 5 spectral window centred on the H emission line. Regions with peak intensities exceeding 3 above the background were selected and H fluxes were measured within circular apertures of radius . We make direct use of the regions identified and characterized in their analysis; the details of the identification and measurement procedures can be found in their paper. The greyscale background in Fig. 10 displays continuum-subtracted H line emission image from MUSE, integrated over the wavelength range 6591.54–6605.29Å. The plot illustrates the projected spatial relationship between the regions and clumps.
@@
We associate the clumps identified in Section 3.4 with the regions based on their proximity in both spatial location and velocity. It turns out that 4 clumps (ID: 1, 2, 4, 5) have associated regions, which are indicated in Table 2. The star formation rate (SFR) for the whole system and each individual region was calculated from the line flux (corrected for extinction using the Balmer decrement technique) of the regions, employing the calibration provided by Kennicutt (1998), which assumes a Salpeter initial mass function (IMF). However, we note that this –SFR conversion may be uncertain in our case, as the assumption of constant (stationary) star formation over timescales of several Myr may not hold. Moreover, at such low SFR levels (especially for ), stochastic sampling of the IMF may lead to an underestimation of the total star formation (see e.g. Boselli et al. 2009, 2018a).
@@
In Fig. 11, we explore the relation between the SFR and mass of the clumps. Literature samples of nearby galaxies and AGC 226178 as a whole are also plotted for comparison. There is a correlation between the mass and SFR, but with substantial scatter. The clumps associated with star formation (blue squares in Fig. 11) in AGC 226178 generally follow the relation observed for nearby galaxies (left panel). The -based star formation efficiency SFR/ of AGC 226178 and its clumps fall at the lower end of the range spanned by nearby galaxies (right panel). This broad consistency of AGC 226178 with normal galaxies supports the view that the -SFR relation does not reflect a causal connection, but rather is a manifestation of star formation being self-regulated by heating (partly contributed by star formation) and radiative cooling of the interstellar medium (e.g. Taylor & Webster 2005).
@@
In Fig. 12, we further explore the structural properties of the clumps and their connection to star formation. The mass is plotted against clump size in the left panel, and the velocity dispersion of the clumps is plotted against size in the right panel. Mass-size relations corresponding to several characteristic surface densities are also plotted for comparison. Particularly, an surface density of is about the observed maximum on (sub)kiloparsec scales of nearby galaxies (e.g. Wong & Blitz 2002). It is thought that at higher surface densities is mostly transformed to molecular hydrogen. A surface density of is about the theoretical critical value for a gas cloud (in protogalactic discs) that can be self-shielded from cosmic background radiation (Taylor & Webster 2005).
@@
In the left panel of Fig. 12, we also overplot the five almost dark clouds (including AGC 226178) from Cannon et al. (2015). There is a tight mass-size correlation for these sources, similar to the well-known tight correlation observed for ordinary galaxies (e.g. Broeils & Rhee 1997; Wang et al. 2016). The sense of the correlation is broadly consistent with nearly constant average surface densities in different systems. The existence of a similarly tight mass-size correlation for clumps and almost dark clouds suggests a self-regulated structural property of interstellar neutral gas, regardless of environment. With that being said, we note that all of the sources plotted in Fig. 12 lie below the line of maximum surface density mentioned above, suggesting that these clouds or clumps are -dominated. In addition, the star formation-free clumps have systematically lower surface densities than the clumps with star formation and other almost dark clouds. The average surface densities of the star formation-free clumps are systematically lower than the theoretical mentioned above. The intracluster radiation field (ignored in the theoretical calculation) is generally much stronger than the cosmic background radiation field. It is therefore quite possible that the clumps without associated star formation are not sufficiently self-shielded, which may inhibit the formation of molecular hydrogen and, consequently, suppress star formation.
@@
In the clump size-velocity dispersion plot shown in the right panel of Fig. 12, there is a significant correlation (the Pearson correlation coefficient = 0.86), which is expected for a hierarchical fragmentation governed by turbulence. The best-fit relation is
@@
The logarithmic slope of 2.77 is much steeper than that ( 0.4-0.5; e.g. Larson 1981; Heyer et al. 2009) of the Larson relations for molecular clouds. A cloud dominated by self-gravity has a logarithmic slope close to 0.5, by assuming a constant surface density. Therefore, our clumps are not expected to be bound by gravity. In addition, we note that the velocity dispersion of clumps associated with star formation is systematically (for a given size) larger than those not associated with star formation. This suggests a significant feedback of star formation on the neighbouring gas.
@@ 4 Discussion
4.1 The connection between AGC 226178 and nearby galaxies
@@
Jones et al. (2022a) uncovered a faint, low signal-to-noise bridge between AGC 226178 and VCC 2034, based on the ALFALFA observation. If the bridge were confirmed, it would be a strong evidence for a physical connection between the two systems. Based on this discovery, Jones et al. (2022a) came to the conclusion that AGC 226178 was formed from a gas stripped from VCC 2034. The unusually high gas-phase metallicity ( 0.5 ) of AGC 226178 indicates that its gas was probably from a galaxy with a much higher stellar mass ( 108.4 ; Jones et al. 2022a) rather than self-enriched. The model from Junais et al. (2022) showed that VCC 2034 may have experienced a ram-pressure stripping event starting 150 Myr ago, losing about of gas, which could comfortably account for the formation of AGC 226178.
@@
Besides the neutral gas, ionized gas emission may also be used to probe the environmental stripping event (Jones et al. 2022b). According to the radiative hydrodynamical simulations of Lee et al. (2022), besides the relatively bright H emitting star-forming regions, there exists diffuse ionized gas in the ram pressure stripped tails. Diffuse H emission from the stripped tails may be detected at surface brightness levels fainter than 6 1038 erg s-1 kpc-2. This is within the reach of the Virgo Environmental Survey Tracing Ionised Gas Emission (VESTIGE) survey (Boselli et al. 2018b), whose 1 surface brightness sensitivity is 3.4 1038 erg s-1 kpc-2 for a distance of 16.5 Mpc (Jones et al. 2022b). However, as discussed in detail by Boselli et al. (2022), this ionized phase may be short-lived. In RPS-affected galaxies, gas is likely stripped as neutral hydrogen and then becomes ionized and eventually heated when mixing with the intracluster medium (ICM). Consequently, diffuse ionized gas is only observed in a handful of late-type galaxies in Virgo, most of which are currently undergoing active ram pressure stripping (Longobardi et al. 2020; Sardaneta et al. 2022). Nevertheless, given the signs of potential environmental interaction in AGC 226178, we examined the VESTIGE narrow-band H image, we did not find any diffuse emission in between and beyond the optical main body of AGC 226178 and VCC 2034 at the current epoch.
@@
4.2 The nature of AGC 226178: A stripped disintegrating gas cloud vs a mini-halo